### v4

# Reexamining Paradigms of End-to-End Data Movement

Abstract.
The pursuit of high-performance data transfer often focuses on raw network bandwidth, where international links of 100 Gbps or higher are frequently considered the primary enabler. While necessary, this network-centric view is incomplete. It equates provisioned link speeds with practical, sustainable data movement capabilities. It is a common observation that lower than desired data rates manifest even on 10 Gbps links and commodity hardware, with higher-speed networks only amplifying their visibility. We investigate six paradigms – from network latency and TCP congestion control to host-side factors such as CPU performance and virtualization – that critically impact data movement workflows. These paradigms represent widely “accepted” engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. We introduce the “Drainage Basin Pattern” conceptual model for reasoning about end-to-end data flow constraints across heterogeneous hardware and software components at varying desired data rates to address the fidelity gap between raw bandwidth and application-level throughput. Our findings are validated through rigorous production-scale deployments, from 10 Gbps links to U.S. DOE ESnet technical evaluations and transcontinental production trials over 100 Gbps operational links. The results demonstrate that principal bottlenecks often reside outside the network core, and that a holistic hardware-software co-design enables consistent, predictable performance for moving data at scale and speed.
End-to-End Data Movement, Co-design, Data Movement Appliances, Burst Buffer, System-level Optimization, Cloud, Cross-Domain Data Flow, Research and Education Networks
$c$s: Computer systems organization Distributed architecturesccs: Computer systems organization Heterogeneous hardwareccs: Networks Network performance analysisccs: Information systems Storage management
Preprint
This is a preprint of a research manuscript currently undergoing peer review.
The work introduces a co-design principle for moving data efficiently and the Drainage Basin Pattern (Fig. 1) conceptual model. Successive implementations have undergone over a decade of validation through production-scale deployments from 10 Gbps to 100 Gbps production links, including U.S. Department of Energy (DOE) ESnet technical evaluations (Kissel and Fang, 2020c, b, a) and historical transcontinental production trials (ICM, 2019).
https://creativecommons.org/licenses/by/4.0/
For the latest version, please see:
https://arxiv.org/abs/2512.15028
Authors’ Contact Information:
Chin Fang, Zettar Inc., USA, fangchin@zettar.com
Timothy Stitt, F. Hoffmann-La Roche Ltd., Switzerland
Michael J. McManus, Intel Corporation (Retired), USA
Toshio Moriya, Institute of Materials Structure Science, KEK, Japan
For academic discussion only.
1. Introduction
The efficient movement of data from edge locations to core data centers and cloud resources is an unsolved foundational challenge for modern data-intensive enterprises and scientific collaborations. High-speed long-distance networks, celebrated for raw bandwidth, are a critical response to this challenge. For example, ESnet6 provides high-capacity transcontinental and international connectivity essential for global scientific collaborations, with speeds from 400 Gbps to 1.2 Tbps (Energy Sciences Network (2025a), ESnet). However, a predominant focus on high raw network bandwidth often obscures a more critical truth: sustainable end-to-end data movement is constrained by the full environment along the data path— including storage, host architectures, software design, and security measures—rather than by the network alone. This holds true from 10 Gbps to 100 Gbps and faster.
The analyses in this paper are based on real-world, production-scale deployments, e.g., transcontinental links, edge-to-core transfers, i.e., operational environment-based. This operational grounding ensures that our insights directly inform system design and deployment practices, rather than being limited to synthetic or small-scale testbeds.
We identify this discrepancy as a “fidelity gap” between theoretical link capacity and actual application-level throughput. This gap manifests as lower than desired data rates, even on 10 Gbps links and modest commodity hardware. Higher-speed networks only amplify its effect. The “Drainage Basin Pattern” conceptual model (Fig. 1) addresses it by shifting the focus from isolated network optimization to systemic co-design. The efficacy of this approach is evidenced by achieving near-line-rate performance on a “high-speed” 100 Gbps transcontinental link (~84 Gbps in a pharmaceutical production environment). It also enables efficient 1–10 Gbps transfers at the edge. Detailed methodology and measurement results appear in Sections 3 and 4, where we analyze how host configurations influence end-to-end throughput across multiple deployment scenarios. In this paper, we define a “fast” network as 10 Gbps up to (but excluding) 100 Gbps, and a “high-speed” network as 100 Gbps and faster.
Figure 1. The experience of moving data for most practitioners is typically limited to the source of the river. This corresponds to end-user activities, such as transferring photos and videos from mobile phones or moving spreadsheets from one folder to another. At such low data rates, operations appear simple. This may explain why efficient data transfer has historically received limited attention, motivating the introduction of the Drainage Basin Pattern.
A conceptual diagram showing the Drainage Basin Pattern for data movement. Illustrates how data flows from edge sources through networks to core data centers, similar to water flowing from streams into rivers to ocean.
Drawing on over a decade of operational experience—spanning extreme scientific workloads at LCLS-II (Thayer and Fang, 2019), sustained petabyte transfers (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet), and international deployments (ICM, 2019)—our work demonstrates that once a link is in place, whether fast or high-speed, host-side configurations become the primary determinants of performance. These architectural principles have been validated through rigorous production-scale deployments, including U.S. DOE ESnet technical evaluations (Kissel and Fang, 2020c, b, a), historical transcontinental trials over 100 Gbps operational links, and high-performance validations on the TaiWan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (2025), TWAREN). Together, this body of evidence confirms that the bottlenecks have shifted from the wide-area fabric to the internal systems architecture of the endpoints.
A key architectural component for achieving predictable performance is the use of high-speed burst buffers (Bhimji et al., 2016) as staging areas, enabling reliable, near-line-rate movement when the end-to-end path is properly co-designed. Our prior work, including the Overall Winner at the Inaugural Supercomputing Asia 2019 Data Mover Challenge (SCA19 DMC) (NSCC, 2019) and official ESnet evaluations (Kissel and Fang, 2020c, b, a), underscores that these factors, not the network, set the true performance limits of data movement, regardless of the network’s nominal speed.
The core contribution of this paper is the identification and validation of six widely held engineering paradigms. They continue to shape expectations, system design, and investment decisions in production end-to-end data movement. Through production-scale deployments and operational trials, we show why they often fail to predict achievable performance across the full edge-to-core spectrum. Reliable and predictable data movement emerges from holistic, co-designed, and tightly integrated hardware-software systems, instead of isolated optimization of individual components. The end-to-end behavior of these systems can be reasoned about using the proposed Drainage Basin Pattern.
The remainder of this paper evaluates the six paradigms through production-scale deployments and examines how the proposed principles generalize to emerging network speeds beyond 100 Gbps.
The six paradigms we re-examine are: (1) the assumed dominance of network latency as the primary performance constraint; (2) the impact of packet loss and the emphasis placed on standard TCP Congestion Control Algorithms (CCA); (3) the presumed necessity of dedicated private lines for high-speed performance verification; (4) the expectation of a direct relationship between increased link bandwidth and application transfer rates; (5) the belief in the essential need for premium CPUs at every endpoint; and (6) the assumed universal utility of virtualization and cloud environments for high-throughput workflows. These paradigms represent widely adopted engineering assumptions that inform system design, procurement decisions, and operational practices in production data movement environments. Throughout this paper, all referenced systems are assumed to run Red Hat Enterprise Linux (RHEL) 9.6 (Red Hat, 2025b) or free rebuilds such as Rocky Linux 9.6 (Rocky Enterprise Software Foundation, 2025).
2. Background / Related Work
Research programs such as the Global Research Platform (GRP) ((2025a), GRP) have successfully promoted the orchestration of high-speed networks and dynamic path provisioning. A recent workshop ((2025b), GRP) and similar community efforts have significantly increased the visibility and understanding of programmable 400 Gbps-class connectivity. However, a predominant focus on the network can lead to treating the data mover software and storage as auxiliary considerations. This work provides a re-balanced perspective by examining data movement as a first-class systems problem (Fang et al., 2016; Fang, 2019).
Rather than revisiting individual network mechanisms, we focus on how storage behavior, host architecture, and data mover design interact with high-speed networks in sustained production workflows. This systems-oriented framing shifts attention from isolated component optimization to end-to-end operational behavior.
2.1. Burst Buffers and Data Staging
We now introduce the burst buffer (Bhimji et al., 2016)—a high-speed intermediate storage layer. Originally pioneered in supercomputing to absorb data for high-performance computing (HPC) nodes at extreme velocities, we generalize this concept for sustained, wide-area data movement.
At the hardware level, burst buffers are built on Non-Volatile Memory express (NVMe) Solid State Drives (SSDs). In practice, this storage is formed by a dedicated software layer—such as a local file system or a scale‑out file store—that provides the underlying storage abstraction.
The burst buffer serves both as a staging area between production storage and the network. A vendor-agnostic data movement appliance design employs a unified software data mover (Zettar zx (Zettar Inc., 2025). See Table 1 for a summary of its features and capabilities) and burst buffers to create a predictable performance envelope that enables sustained near-line-rate data movement between burst buffers (fast ($100~\mathrm{Gbps}$) and high-speed ($\geq 100~\mathrm{Gbps}$) WAN links)
In this architecture, the burst buffer serves both as a fast storage tier and as a deliberate decoupling mechanism. It isolates the data mover from production storage services, which are often optimized for capacity or ease of use, rather than throughput or latency. This decoupling allows each layer to operate within its effective performance regime while staging coordinates data movement across mismatched behaviors.
Data staging, often defined simply as the movement of data between production storage and burst buffers, is a critical coordinating process. This operation must be straightforward, predictable, and highly efficient, as any delay in staging fundamentally negates the performance benefits of burst buffering.
The need for staging arises because production storage systems are frequently constrained by throughput, latency, and protocol limitations, making them incapable of feeding high-speed WAN links directly.
This role becomes increasingly critical at aggregation points, where multiple modest-rate flows converge and expose the mismatch between distributed edge behavior and centralized storage services. Without effective staging, such convergence shifts the bottleneck from the network to storage, negating the benefits of high-speed connectivity.
To meet the need, the unified data mover zx was developed as the foundation for the holistic co-design principle central to this work. Its architecture integrates all necessary functions— including over-distance transfer and internal data staging (between production storage and burst buffers), supporting diverse protocols (file and object), and handling both bulk and streaming transfers. This single, concurrent, and scale-out data mover manages the complete data placement workflow, from source storage through transit to destination storage. Burst Buffers and Data Staging are thus a foundational element of our Co-design Engineering Principle.
2.2. Bulk and Streaming Transfers
The data movement workloads supported by zx fall into two primary types:
  • Bulk Transfer: the movement of a static, pre-existing dataset where the complete data is at rest in the source storage before the transfer initiates. This is typical for migrating archives, databases, or completed experimental runs.
  • Streaming Transfer: the movement of a dynamic, actively growing dataset where data is transferred concurrently as it is being generated and written to the source storage. This type is essential for real-time workloads, such as data acquisition from scientific instruments like LCLS-II (Thayer and Fang, 2019) and large-scale bioscience imaging operations (e.g., Cryo-EM, virtual staining) that, due to their inherent data rates, will require such concurrent, line-rate data movement to achieve efficient production workflows.
This distinction is critical because streaming transfers demand both high storage throughput and low storage latency, stressing storage subsystems, memory hierarchies, and scheduling mechanisms in ways that differ fundamentally from bulk transfer. As a result, streaming workloads require architectures designed for sustained concurrency and flow.
Traditional High-Performance Computing (HPC) environments commonly integrate data movement into centralized batch-scheduled workflows, in which transfers are initiated, throttled, and terminated by external resource managers. While effective for coarse-grained compute allocation, this model introduces non-deterministic start times, queueing delay, and run-to-run jitter into the data path. These effects are not incidental: by treating data movement as a schedulable job rather than a continuously executing system service, centralized orchestration disrupts the temporal continuity required for sustained, near-line-rate streaming.
The architecture presented in this work eliminates this source of variability by removing centralized orchestration from the data path entirely. Data movement is executed as a decentralized, peer-to-peer service, coordinated implicitly through asynchronous buffer state rather than explicit global scheduling. In this model, transfer cadence emerges from the zx-managed interaction among co-designed hardware resources, the host operating system, and local burst buffers, rather than from external queuing policies. As a result, throughput stability is bounded by physical system limits instead of orchestration-induced control latency, enabling predictable performance across heterogeneous deployment environments.
This design resolves the critical mismatch that occurs when high-rate network ingress meets latency-bound production storage—a problem that lies beyond the expertise of most operational IT staff. It does so without requiring application-level awareness or manual tuning.
2.3. Integrated Appliances vs Software-Centric Approaches
The network-centric approach discussed previously, while valuable, does not fully address the processes required for achieving target data rates reliably in production. Empirical evidence, including our own experience, also indicates that providing software alone—even with proven performance in validation trials (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; ICM, 2019; NSCC, 2019)—can create significant operational challenges. Integration, tuning, and maintenance require deep, multi-disciplinary expertise, presenting a substantial total cost of ownership that many organizations are not equipped to sustain.
The resulting variability makes performance unpredictable and hard to reproduce across different organizations and environments.
This paper demonstrates that the end-to-end data movement problem, despite its variations across workflows, can be addressed through a singular co-design engineering principle. This work builds on extensive prior advances in high-speed networking and host tuning, and focuses on architectural principles that make such expertise deployable and repeatable across production environments. It is this principle that forms the basis for our reexamination of the six paradigms, providing the holistic framework that reveals why these isolated beliefs are insufficient and how to do it right. We present a unified appliance architecture that embodies this principle across the entire data movement workflow spectrum. This co-design principle is applied across a range of hardware form-factors, from compact, low-power nodes and specialized Data Processing Units (DPUs) to certified enterprise servers from tier-1 vendors, supporting data rates from 1 Gbps up to 100 Gbps and beyond. These appliances encapsulate the necessary hardware/software co-design and tuning that was previously a manual, expert-driven process.
The overall approach is grounded in technology developed over the past decade to meet the ambitious data movement requirements of the premier United States (U.S.) Department of Energy (DOE) Exascale Computing Preparation Project, LCLS-II (Thayer and Fang, 2019). The core challenges and performance targets defined by the project directly motivated the architecture. Furthermore, the ongoing dialogue with LCLS-II and other tier-1 scientific entities continues to advance and validate the evolution of this holistic approach to data movement systems. The core software’s performance has been independently validated on long-distance 100 Gbps links, encompassing petabyte-scale transfers with ESnet (Energy Sciences Network (2017), ESnet), (Energy Sciences Network (2018), ESnet), and (Kissel and Fang, 2020c), SCA19 DMC (NSCC, 2019), and transcontinental production trials (ICM, 2019).
The outcomes documented by independent institutions, e.g., (Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; NSCC, 2019) are reproducible; nevertheless, our experience shows that operational reproducibility depends critically on the availability of pre-validated system configurations rather than ad-hoc assembly at each site. This co-design principle has been implemented in commercially available appliances, with implementations ranging from regular server-based units from tier-1 vendors such as Hewlett Packard Enterprise (HPE) to micro servers (Fig. 3). This demonstrates that reliable, efficient data movement is most effectively achieved through complete systems engineering (Kissel and Fang, 2020b, a). The underlying principle is akin to a high-performance electric vehicle (EV): engineered for extreme speed and efficiency, yet perfectly capable of simple, reliable operation for everyday tasks. Just as an EV is not merely a carriage with an electric motor swapped in, a purposely-built data movement appliance is not simply a server with powerful CPUs, NVMe SSDs, and high-bandwidth NICs; the hardware Bill of Materials (BOM) is meticulously formed to maximize operational and cost efficiency. The components are orchestrated via software to eliminate the non-deterministic bottlenecks inherent in general-purpose computing. In passing we will note that attempts to introduce legacy, non-optimized software data movers on such a deterministic platform result in performance regressions, negating the advantages of the co-design principle.
Despite the recognized challenges of end-to-end data movement, the prevailing industry and research approach continues to favor complex tuning methodologies and reliance on prohibitively expensive hardware. Such factors are particularly acute at the network edge — the "headwaters" of the “Drainage Basin Pattern”, Fig.1, where resource-constrained environments such as hospitals, clinics, and remote laboratories typically utilize 1–10 Gbps links. Deploying reliable, high-efficiency data transfer in these settings becomes cost-prohibitive when conventional solutions require specialized IT staff and expensive hardware.
The architecture proposed in this paper directly addresses these barriers by validating a different principle: that architectural co-design can inherently reduce both operational complexity and capital expenditure across the entire data spectrum. This principle is practically realized in a mini-appliance designed for resource constrained edge sites. The hardware costs approximately \$2,000 (Fig. 3 Mini Appliance), including high-performance NVMe burst buffers. This demonstrates that high-efficiency 1–10 Gbps data transfer is achievable simply and affordably, without complex system integration or the utilization of inappropriate virtualization options (Section 3.6). This cost efficiency and simplicity are a direct consequence of the holistic approach detailed in the following sections.
3. Six Common Paradigms
3.1. The Latency Killer Paradigm
A prevalent belief in data-intensive computing is that network latency is the ultimate data movement killer. While this thinking is common among many experienced IT professionals, our empirical testing and architectural analysis motivates the re-examination of this paradigm. The perceived severity of latency is fundamentally governed by TCP's Bandwidth-Delay Product (BDP), which dictates the congestion window required for full bandwidth utilization.
This paradigm can present significant challenges for distributed, data-intensive engineering and scientific endeavors, an impact particularly relevant in the biopharma and life sciences sector—the top global data generator, with data volume projected to grow at a Compound Annual Growth Rate (CAGR) of 36% through 2025 (RBC Capital Markets, 2025). The challenges of moving data at this scale are a primary concern for industry leaders, including organizations such as tier-1 biopharma businesses and enterprise computing providers like HPE. Inefficient data movement, if not properly addressed, can negatively impact drug discovery, precision medicine development, and research timelines. However, as we demonstrate, these challenges can be significantly mitigated through architectural and software approaches that reduce sensitivity to latency.
For years, established resources such as the U.S. DOE ESnet's fasterdata knowledge base (Energy Sciences Network (2025b), ESnet) have documented methods to mitigate latency's impact on data rates. While ESnet provides valuable guidance as a network service provider, a comprehensive solution requires consideration of the entire environment, not just the network—a point we expand in Section 3.3. The reproducible test results summarized in Fig. 2 and the appliance design in Fig. 3 together exemplify this holistic approach. See Table 2 for Ethernet adapter ring buffer settings and Table 3 for tuned Linux kernel parameters.
The test utilized iperf3, a benchmark tool maintained by the U.S. DOE ESnet (Energy Sciences Network (2025g), ESnet). Beginning with version 3.16 (released November 30, 2023 (Energy Sciences Network (2023), ESnet)), the tool was re-architected to be genuinely multithreaded, making it significantly easier to tune for high-latency, high-speed links (e.g., ~160 Gbps memory-to-memory over a 200 Gbps ESnet testbed data path) (Energy Sciences Network (2025d), ESnet).
Section 3.3 provides descriptions of the testbed design. Among many uses of the testbed, it enables the testing of different latency values automatically (Fig. 1). Such a testbed was first constructed in the former Intel Swindon lab, in Swindon, United Kingdom (U.K.). Note that round-trip time (RTT), as reported by tools like ping (man7.org, 2025), is approximately 2 × one-way latency.
For reproducibility, the employed HPE server-based appliance, including network interface card (NIC) settings, tuned Linux kernel parameters, and simulated latency values are described in Tables 2-4. More related materials are publicly available in a GitHub repository (Fang, 2025).
Figure 2. iperf3 latency sweep results obtained using two HPE DL380 Gen 11 server-based appliances (Fig. 7) on a latency simulation-capable testbed (max network speed 100 Gbps) established in the former Intel Swindon Lab, Swindon, U.K., in 2024. While default kernel networking settings (OOTB) show severe performance degradation under high latency, proper kernel tuning substantially reduces this penalty.
A line graph showing iperf3 latency sweep results comparing default vs. tuned kernel settings on a 100 Gbps connection with simulated latencies from 10ms to 100ms.
Table 1. Key functional capabilities and core features of the Zettar zx data movement software.
Capability Notes
Support bulk and streaming transfers Utilizes TCP for transport
Sustained performance independent of Encryption, compression, latency and file sizes
Cluster Architecture Linear scaling capabilities
Single Application Supports File, object, locally and over distance, on-prem and cloud
Symmetric Concurrent Send/Receive rsync and bbcp alike
Embeddable on specialized hardware Support for NVIDIA and Intel DPUs
Quality of Service (QoS) Built-in support for traffic prioritization
Integration with typical IAM Active Directory (AD) and OpenID Connect (OIDC) support
Table 2. rx and tx ring buffer values for Intel® Ethernet Network Adapter E810-2CQDA2.
Parameter Value
rx_value 8160
tx_value 8160
Figure 3. Bill of Materials (BOM) and component details for the Core (HPE DL380 Gen 11) and Mini (Minisforum MS A2) appliances, demonstrating the vendor- and form-factor-agnostic unified data movement appliance design which achieves consistent performance across implementations.
A comparison table showing Bill of Materials and component details for Core (HPE DL380) and Mini (Minisforum) data movement appliances.
Table 3. Tuned Linux kernel parameters. This specific set was empirically determined to be optimal for the zx data mover’s concurrency model, whereas other data movers may require more extensive tuning. For example, see ESnet fasterdata, Linux Tuning (Energy Sciences Network (2025e), ESnet).
Parameter Value
net.core.rmem_max 2147483647
net.core.wmem_max 2147483647
net.ipv4.tcp_rmem 4096 67108864 1073741824
net.ipv4.tcp_wmem 4096 67108864 1073741824
net.ipv4.tcp_mtu_probing 1
net.core.default_qdisc fq_codel
net.ipv4.tcp_congestion_control cubic
net.core.netdev_max_backlog 8192
Table 4. Some common latency ranges. 10 ms, 50 ms, and 100 ms were used for latency simulation.
Distance category Latency range (ms)
Metropolitan (same city or nearby) 1 – 10
Interstate (within the same country) 10 – 50
Cross Continent 50 – 100
Resources such as the ESnet fasterdata knowledge base provide invaluable guidance. However, practitioners sometimes apply published tuning parameters as static prescriptions rather than adaptable guidelines. This approach can lead to suboptimal performance when it doesn't account for interactions between specific data mover software, hardware stacks, and application workloads. Effective high-performance data movement requires understanding and adaptation beyond copy and paste. Note that the results in Fig. 2 and the modest parameter set in Tables 2-3 demonstrate how an architectural approach also can reduce complexity and potential for misconfiguration.
3.2. Universal Packet Loss and TCP CCAs
This section provides a systematic examination of two interconnected paradigms of end-to-end data movement: the assumption of universal packet loss and the impact of TCP CCA selection. We begin by quantifying the data volumes supported by common transfer rates, then introduce a conceptual model—the “Drainage Basin Pattern”—that frames the entire spectrum of data movement workflows. Subsequently, we present empirical evidence that challenges the two prior assumptions in high-speed environments. A real-world case study, drawn from a Wall Street Journal report, illustrates the practical consequences of these paradigms. To clarify the scale of these workflows, Table 5 quantifies the daily data volume achievable at common network speeds. Finally, we discuss why conflating different classes of workflows often leads to suboptimal outcomes.
Table 5. Daily data volume achievable at three common network speeds.
Data rate (Gbps) Data (TB)/day Notes
1 10 5G speed often $\leq$ 0.5 Gbps
10 100 High-performance edge transfer
100 1000 i.e. 1 PetaByte
While the end-to-end movement of data often appears transparent to application users, the underlying mechanisms are complex. The “Drainage Basin Pattern” illustrates this full spectrum of data movement workflows, revealing considerations that may not be immediately apparent.
While most users experience only the “source of the river,” where consumer-grade networks may indeed have packet loss, the “high-speed backbone”—the deep main river channel—is an entirely different environment. This recognition is based on extensive experience since 2015 with 100 Gbps and faster connections, gained through collaboration with the U.S. DOE ESnet and various tier-1 organizations. Consequently, our observations confirm that in the well-engineered 100 Gbps+ Research and Education (R&E) networks discussed here—principally ESnet and Internet2 (Internet2, 2025)—packet loss is effectively negligible during normal operation. This observed reliability is also a key characteristic of peer R&E networks such as GÉANT (GEANT Association, 2025), AARNet (AARNet, 2025), and SingAREN (SingAREN, 2025).
3.2.1. The Packet Loss Paradigm
The nature of data transmission at high speeds differs fundamentally in both engineering and measurement from casual, end-user level data transfers. For high-speed networks, "packets" often vary in size and are transmitted at rates where bit-level impairments dominate over per-packet events. For example, for nodes connected to 10 Gbps and faster networks, it is common to employ "jumbo frame" Maximum Transmission Unit (MTU) of 9,014 or 9,018 bytes, rather than a constant 1,500-byte frame. Consequently, high-speed network device vendors typically specify performance using Bit Error Rates (BER) (National Institute of Standards and Technology, 2025) rather than packet loss. While operators historically reported BERs more explicitly, modern practitioners typically must extract such information indirectly via monitoring infrastructures such as the perfSONAR lookup service (perfSONAR, 2025).
To place the current argument in context, we cite a 2009 ESnet reference to show the historical lower bounds of key performance metrics. Subsequently, we present a practical 2018 case study to demonstrate that packet loss is not the dominant constraint for high-volume, high-speed data transfer. ESnet's Services and Service Level Descriptions (SLD) (Appendix A in (Energy Sciences Network (2009), ESnet)) specifies guaranteed network performance via its Loss Thresholds Table. This table typically lists a Frame Loss Rate (FLR) range of 10-7 to 10-10 and includes a formula for converting the FLR to BER. Furthermore, an accompanying ESnet Fasterdata Knowledge Base (KB) article on Packet Loss provides an approximate translation of BER to end-to-end packet loss, citing a typical scenario of 1 packet out of 22,000 packets, or 0.0046%. Converting this packet loss ratio to a BER illustrates the exceptional quality of these networks:
  • Packet Loss Rate = 0.0046% ($4.6\times 10^{-5}$)
  • Assuming a standard 1500-byte frame: BER $\approx(4.6\times 10^{-5})/(1500~\text{bytes/frame}\times 8~\text{bits/byte})\approx 4\times 10^{-9}$
This result—a BER of $4\times 10^{-9}$, meaning roughly 4 erroneous bits in every billion—confirms that packet loss is negligible in well-engineered backbones. Note that using jumbo frames would further reduce the implied BER.
The operational lower bound on error rates is corroborated by practical, large-scale experience. In September 2018, using the setup shown on page 19, Appendix 6.1 of (Kissel and Fang, 2020c) (see also Fig. 12), a production trial run was carried out for 1 PB transfer in 29 hours over a 5,000-mile 100 Gbps WAN loop. The setup was designed and implemented at SLAC National Accelerator Laboratory, using equipment that, at the time of deployment, was three years past its initial deployment. The trial was reported in (Energy Sciences Network (2018), ESnet). Of the hardware-limited 80 Gbps bandwidth, with full encryption and checksumming, an average utilization of 76.63 Gbps was achieved. This result directly contradicts the notion that packet loss is the dominant limiter in modern, well-engineered wide-area backbones.
3.2.2. The TCP CCA Paradigm
The empirical evidence presented previously established the irrelevance of packet loss on well-engineered network backbones. Next, we examine the related paradigm that the choice of TCP CCA is critical for performance. The prevalence of this view is evident from the abstract of Google Networking Research's publication "BBR: Congestion-Based Congestion Control" (Cardwell et al., 2016), which introduced the BBR algorithm (Bottleneck Bandwidth and Round-trip Propagation Time). The abstract states: “Physics and climate researchers need to exchange petabytes of data with global collaborators but find their carefully engineered multi-Gbps infrastructure often delivers at only a few Mbps over intercontinental distances.”—a claim that implicitly attributes poor performance to TCP behavior rather than end-system or architectural factors.
It is a telling contrast that during the very period this paradigm was being articulated, a production reality was demonstrated. ESnet publicly reported our 1 PB transfer, which achieved 76.63 Gbps average utilization (Energy Sciences Network (2018), ESnet). This result was accomplished not with a novel CCA, but with the default CUBIC in the standard CentOS 7.5 distribution, which did not yet include BBR. This demonstrates that the fundamental bottleneck was not the CCA, but the lack of a holistically co-designed system.
Before proceeding, we establish the context regarding high-speed TCP CCAs by citing relevant Internet Engineering Task Force (IETF) Request for Comments (RFCs) and referring to prior work (Energy Sciences Network (2018), ESnet). Starting in 2018, the suitability of CUBIC for Fast Long-Distance Networks was already documented by RFC8312 (Rhee et al., 2018). Coincidentally, earlier work (Energy Sciences Network (2018), ESnet) described a production trial conducted on a 5000-mile, 100 Gbps loop, achieving over 96% bandwidth utilization—practically, line rate. The systems used in this trial ran CentOS 7.5-1804, which utilized CUBIC as its default CCA. We note that RFC9438 (Xu et al., 2023), titled “CUBIC for Fast and Long-Distance Networks”, has since obsoleted RFC8312. More recently, RHEL 9.6 and its free rebuilds now include BBR as the default CCA, as demonstrated by the simple command output presented below (command line output wrapped to fit in a single column):
\$ sysctl net.ipv4.tcp_available_congestion_control
net.ipv4.tcp_available_congestion_control = reno cubic bbr
It is noteworthy that while Google researchers introduced the model-based BBR CCA in 2018—the same year CUBIC was already the default in widely used Linux operating system distributions such as CentOS 7—BBR has since progressed to version 3 (Google, 2025). However, the BBR implementation that ships with RHEL 9.6 remains version 1, which is notably more aggressive than later versions. This distinction is relevant because BBRv1’s aggressiveness should, if anything, favor higher throughput in uncongested environments. Here, we compared the performance of BBR, CUBIC, and Reno (a classical TCP CCA algorithm that emerged in the 1990s) and provided the results in Figs 4-6. The following definitions and methodology are established before presenting the three figures.
  • Hyperscale Data Set: A synthetic test data set containing files of a uniform size, where the total number of files is 220 (1,048,576) or the aggregate size is 1 TiB, or both.
  • Data transfer sweep:
    • Bulk Transfer Sweep: File sizes range from 1 KiB to 1 TiB, with sizes incremented by powers of two, resulting in 31 distinct datasets.
    • Streaming transfer sweep: File size ranges from 4 MiB to 1 TiB; sizes are incremented in power of two, so there are 19 datasets
Each sweep iteration processed datasets sequentially from the smallest to the largest file size. This process was repeated for multiple iterations to gather statistically stable results.
Note that statistical analysis for each complete sweep (bulk and streaming) included the calculation of the mean, median, and standard deviation. The mean values are plotted automatically using a custom gnuplot command file (Gnuplot Development Team, 2024). See the GitHub repo (Fang, 2025) for tables containing raw mean, median, and standard deviation values.
Figure 4. A bulk transfer sweep leveraging kTLS (kernel TLS) offload in RHEL 9.6 (Red Hat, 2025a) to evaluate the three default congestion control algorithms (CCAs). BBRv1 did not demonstrate a clear performance benefit over CUBIC or Reno. Incidentally, kTLS did not yield any transfer rate improvement in this case.
Performance graph showing bulk transfer rates vs. file sizes for BBR, CUBIC, and Reno congestion control algorithms with kTLS offload.
Based on the demonstrated functional equivalence of high-speed performance, we conclude that the choice among competing CCAs is not a primary concern for high-volume, high-throughput data transfer. CUBIC is empirically validated, using the appliances in Fig. 2, as a safe and effective default (Figs. 5-6).
Figure 5. Subsequent bulk transfer sweeps were performed without kTLS due to the prior degradation. Using the default RHEL 9.6 CCAs, BBRv1 and CUBIC exhibited identical throughput across 4 MiB–1 TiB. Reno showed degradation consistent with older congestion control paradigms.
Performance graph showing bulk transfer rates vs. file sizes for BBR, CUBIC, and Reno congestion control algorithms without kTLS.
Figure 6. As a unified software data mover (Zettar Inc., 2025), zx has built-in streaming capability, motivated by (Thayer and Fang, 2019). BBRv1 shows no advantage over CUBIC, even for streaming transfers. kTLS was not used.
Performance graph showing streaming transfer rates vs. file sizes for BBR, CUBIC, and Reno congestion control algorithms.
It is worth noting that Reno, characterized by slower window size adjustments than both CUBIC and BBRv1, is highly susceptible to the non-offloaded userspace overheads of TLS. The implementation of kTLS effectively mitigates this susceptibility, allowing Reno to achieve throughput parity with the other CCAs by removing the userspace processing bottleneck (Fig. 4).
3.2.3. Practical Implications and Real-World Consequences
The engineering realities discussed thus far have significant, though often invisible, operational and economic impacts. The tangible consequences of suboptimal data transfer are starkly illustrated by recent extreme measures reported by the Wall Street Journal (WSJ) (WSJ Staff, 2025). Given the exponential data growth worldwide, the mobility of multiple petabytes (PBs) is now a common requirement.
A June 12th article (WSJ Staff, 2025) titled “Chinese AI Companies Dodge U.S. Chip Curbs by Flying Suitcases of Hard Drives Abroad” detailed how four engineers transported 4.8 PB of training data by flying sixty hard drives from Beijing, China to Kuala Lumpur, Malaysia. This volume of data, sufficient for training large language models (LLMs) such as OpenAI’s ChatGPT, highlights the severe real-world constraints imposed by inefficient electronic data transfer.
A related article from Tom's Hardware (Morales, 2025) clarified that this physical transport was a “meticulously planned operation and took several months of preparation. In sharp contrast, the performance results illustrated in Figs. 4-6, combined with the entries in Table 5, show that transferring this volume of data electronically, using a 100 Gbps link, would require at most one week of sustained transfer.
This dichotomy recalls the hierarchy of solutions presented in Figs. 2-3, whose importance can be illustrated by a simple metaphor: providing water. Watering plants on a balcony requires a watering can. Watering a garden demands a hose. However, supplying a metropolitan area from a reservoir necessitates a massively engineered system, e.g., the Hetch Hetchy Aqueduct (Bay Area Water Supply and Conservation Agency (2025), BAWSCA).
This hierarchy directly parallels data movement. Transferring data at rates of ~1 Gbps is akin to using a watering can. Rates of ~10 Gbps correspond to the garden hose. But when approaching 100 Gbps and beyond, the endeavor demands the engineering equivalence of a metropolitan water transport system—a holistic, co-designed infrastructure where ad-hoc solutions fail.
3.3. Dedicated Private Lines Are Essential for High-Speed Testing
A common paradigm asserts that validating high-speed data transfer requires a dedicated, high-bandwidth WAN link. This belief creates a significant barrier, as such links are costly and complex to provision, requiring specialists to operate well. We confronted this directly in 2023 while developing data movement appliances with Intel Corp. Neither organization had a 100 Gbps WAN, raising a critical question: how could we rigorously validate performance in the absence of a production-grade link?
From 2015-2019, the zx R&D was carried out on a testbed at SLAC National Accelerator Laboratory, using a 5,000-mile, 100 Gbps loop provisioned by ESnet. After 2019, collaboration with ESnet continued (Kissel and Fang, 2020c). Nevertheless, accessing their 100G Software Defined Network (SDN) Testbed required a formal application process (Energy Sciences Network (2025f), ESnet), hindering agile development. This need for a fully controlled, on-demand test environment led us to evaluate commercial WAN simulators, which we found to be prohibitively expensive and functionally limited.
A turning point was the discovery of a 2012 technical report, “Validating Linux Network Emulation” (Kissel and Swany, 2012), which outlined a method using Linux's built-in traffic control (tc (The Linux man-pages project, 2025b)) and network emulation (tc-netem (The Linux man-pages project, 2025a)) tools. Recognizing the potential of this software-based approach to create a high-fidelity, 100 Gbps capable testbed at a fraction of the cost, we implemented an enhanced version. This version extended functionality and automation beyond the original method. The resulting testbed (Fig. 7) —built at the former Intel Swindon Lab in collaboration with Intel and HPE—consists of two HPE DL380 Gen 11 servers (configured as the appliances under test) connected via a switch to two Intel "latency servers." These servers use tc and tc-netem to impose precise delay, accurately simulating transcontinental links in an automated, remotely manageable setup.
Figure 7. Intel Corp. arranged its former Swindon Lab in Swindon, U.K., to collaborate with Zettar using an enhanced version of the approach created by Dr. Ezra Kissel. The essential components are labeled explicitly.
The topology of the testbed setup at Intel Swindon Lab showing two HPE servers connected via latency simulation servers.
Figure 8. Bulk transfer sweeps vs three simulated latency values: 10 ms, 50 ms, and 100 ms, corresponding respectively to Los Angeles to San Francisco, California, U.S.; Singapore to Alaska, U.S.; and Singapore to Atlanta, Georgia, U.S.
Graph showing streaming transfer performance vs. simulated latency values (10ms, 50ms, 100ms) on the 100 Gbps-capable testbed.
The results, shown in Fig. 8 (bulk transfers) and Fig. 9 (streaming transfers) demonstrate the testbed's capability to sustain high data rates under simulated latencies of 10ms, 50ms, and 100ms. The critical validation comes from comparing the testbed's performance across this simulated latency range with the results from the production 100 Gbps link, which operates at an approximate latency of 74ms (Fig. 5-6). The achieved data rates and performance profiles are strikingly similar, demonstrating the testbed's predictive reliability across the critical WAN latency range.
The true significance of this Linux-based emulation approach is that it transforms the high-speed development environment from a logistical burden into a strategic, versatile engineering platform. This approach provides a compact, low Total Cost of Ownership (TCO) platform capable of reliably predicting performance profiles, making it a useful engineering tool for development and optimization across numerous critical functions:
Figure 9. Streaming transfer sweeps vs three simulated latency values, 10ms, 50ms, and 100ms. Note that the data rate levels are quite close to their counterparts from the bulk transfer sweeps shown in Fig. 6.
Graph showing streaming transfer performance vs. simulated latency values (10ms, 50ms, 100ms) on the aforementioned 100 Gbps-capable testbed.
  • Systems Focus: Unlike commercial WAN simulators or network testbeds that isolate network performance (often using Random Access Memory (RAM)-to-RAM methods), this platform is uniquely suited to measure true burst-buffer to burst-buffer data transfer rates. This is essential because, as established in Section 3.1, persistent I/O is the dominant constraint for petascale transfers.
  • Agile Development: The platform supports rapid software iteration through automated regression testing for Continuous Integration/Continuous Deployment (CI/CD) and accelerated development of new components, such as AI Model Context Protocol (MCP) with guardrails for connecting zx to a natural language input for improved usability, without risking a live production network.
  • Operational Utility: Its reproducibility enables critical regulation validation for compliance-driven sectors, supports WAN planning through latency simulation, facilitates training/education for new systems engineers, and provides an efficient demo showcase environment.
While this method validates core performance metrics under controlled conditions, it is important to note that it does not replace the necessity of an active SDN testbed for validating complex control-plane, security, or multi-flow operational scenarios typical of a fully instrumented network like ESnet. However, for core high-throughput systems performance engineering, the software-defined WAN emulation offers a cost-effective, high-fidelity alternative.
The full testbed benefits and implementation details, together with expanding it to speed $\geq 100$ Gbps, warrant a separate, in-depth tutorial. See also Section 7.
3.4. Increasing network bandwidth increases transfer rates
In essence, this paradigm fails because the rest of the infrastructure—primarily storage and compute, combined with the parallelism and concurrency of the software data mover—must be in balance with the available network bandwidth. A chain is only as strong as its weakest link; once the network ceases to be that link, introducing more bandwidth yields no benefit.
Foremost among these limiting factors, and often underestimated, is storage I/O. Once network bandwidth exceeds the system's ability to read from or write to persistent storage, further increases are futile. The bottleneck most frequently occurs during write operations, as virtually all storage media deliver lower write than read performance.
Figure 10. For production storage to support moving data at scale and speed well, it must have high enough throughput and low latency.
A diagram illustrating storage throughput and latency requirements needed to support data transfer at speed and scale.
Other system-level factors also contribute: the processing of TCP/IP stacks, filesystem metadata operations, and encryption/checksumming all consume finite CPU cycles that could otherwise be allocated to sustaining storage I/O. As indicated in Fig. 10, to attain high and sustainable data transfer performance, the storage service must meet two criteria: 1) high enough storage Input/Output Operations Per Second (IOPs)/throughput and 2) low latency. Both are elaborated below.
In Fig. 10, the term “high enough” means the overall aggregated storage IOPs/throughput > the target transfer rate. “Low latency” is particularly crucial because, for highly concurrent data movers, storage latency effectively sabotages concurrency—a scenario often associated with random I/O. What appears as a software bottleneck is frequently a storage latency problem in disguise.
Furthermore, as shown in Fig. 1, without a system (including kernel tuning) holistically co-designed to match the data mover's profile, network latency will cripple performance regardless of the raw bandwidth available.
The preceding analysis and evidence support a re-evaluation of this paradigm.
3.5. Powerful CPUs are essential for high transfer rates
A common consensus in high-performance computing is that powerful CPUs are essential for high transfer rates. During 2023, while discussing with our industry collaborators regarding CPU selection for data movement appliances, the prevailing feedback was to select high-end, high-core-count models. We held reservations about this approach. Furthermore, we also had skepticism about suggestions to utilize CPUs with embedded hardware acceleration such as Quick Assist Technology (QAT) (Intel, 2025b), based on the observation that such acceleration can tie software to specific vendor hardware and drivers, increasing complexity and potential "software bloat."
This skepticism was grounded in years of deployment experience, which revealed that zx consistently achieves high performance without requiring a high core count; typically, 12-24 cores are sufficient. While resources like ESnet's fasterdata recommend high clock rates for good encryption performance (Energy Sciences Network (2025c), ESnet), our observations suggested an alternative path. We hypothesized that fewer cores with CPU-built-in encryption acceleration and moderate clock rates would reduce context switches (improving software efficiency) and lower energy consumption, thereby reducing the TCO. This led to the selection of the Intel Xeon 5418N, a mid-range model (Intel, 2025a).
In 2024, more extensive testing with the two HPE DL380 Gen11 server-based appliances confirmed that even with full encryption, QAT was unnecessary. The CPU's native instruction set extensions were sufficient. The results shown in Figs. 4-6, 8-9 were all achieved without hardware acceleration and using this modest CPU.
As such, we can conclude that CPU raw computing power matters, but software efficiency and storage architecture matter more. The right software makes adequate CPUs perform exceptionally; the wrong software cannot be saved by the most powerful CPUs.
3.6. Virtualization and cloud are universally useful
We recently received an inquiry about evaluating zx for a site in South Africa needing to transfer data to Europe at higher rates. The site's proposal to use a virtual machine (VM) immediately raised concerns, as it revealed a common misconception: that virtualized environments can deliver the performance required for high-speed data movement.
Virtualization technologies have been a mainstay for over two decades (Red Hat, 2025c). Nevertheless, a critical determinant of their performance for intensive data movement workloads is the level of system control accessible to the user. Table 6 summarizes the scope of Linux kernel parameters across virtual machines and containers.
Table 6. Linux Kernel Parameter Scope: Virtual Machine vs. Container
Location Platform Scope
Inside VM VM (e.g. KVM) Affects only the guest OS kernel.
Container Docker/Podman Affects namespaced subset; no host change.
Privileged Docker/Podman Can change host settings, affecting all.
On Host Physical host Affects physical server and all VMs and containers.
This hierarchy of control directly leads to three critical performance implications for high-speed data transfer:
  1. 1.
    Interrupt Overhead: Even with Single Root I/O Virtualization (SR-IOV) (NVIDIA, 2025a), interrupt handling in VMs incurs measurable overhead due to interception and remapping by the hypervisor.
  2. 2.
    I/O Performance Degradation: High-speed I/O operations, which generate a high volume of interrupts, consequently achieve lower performance in virtualized environments.
  3. 3.
    Host-Level Dependency: If the host kernel is not tuned to mitigate latency (as established in Section 3.1), adjustments within the guest OS are largely ineffective.
Achieving consistent, latency-insensitive data transfer performance within a VM is intrinsically constrained by the architecture. Cloud providers inherit these limitations. Coordinating the kernel parameter adjustments necessary for maximum VM performance is application-dependent (see Table 3) and practically infeasible for a heterogeneous customer base.
3.6.1. The Architectural Cost of Cloud Abstraction
While virtualization and cloud abstraction deliver significant benefits in manageability and elasticity, they impose substantial performance penalties on data movement. The canonical cloud data path—encumbered by layers of overheads such as hypervisor, virtualized storage, and HTTP/REST APIs—incurs a substantial performance inefficiency compared to a tuned bare-metal environment. This inefficiency is empirically observed in our tests to routinely reach 30–50%. For instance, on 10 Gbps Elastic Compute Cloud (EC2) instances in 2019, we achieved a maximum of ~6 Gbps per VM, a 40% loss from line rate. More recently, as detailed in Fig. 11, the native cloud tooling (aws-cli) performed at a small fraction of the available capacity, while our co-designed appliance, even when partially firewalled, consistently attained a significantly higher percentage of the physical link's potential.
Standard cloud workarounds, such as Amazon Web Services (AWS) multipart uploads (Amazon Web Services, 2025), fragment a simple file transfer into a multi-step coordination process—effectively a highly fragmented workflow (Acharya and Sirinterlikci, 2010) for data movement.
3.6.2. A Critical Examination of Cloud Provider Metrics
The prevailing paradigm of general-purpose cloud data paths, regardless of provider, introduces inherent architectural conflicts for high-performance data movement workflows. This challenge becomes acute at the petascale level. Analysis of major U.S. cloud providers reveals a focus on theoretical network bandwidth, often decoupled from storage I/O and end-to-end performance. The info in Tables 7-9, extracted from official provider documentation, illustrates this focus.
The Disconnect Between Theoretical Bandwidth and Practical Throughput
Maximum network bandwidth figures—such as 1,000 Gbps or 400 Gbps —are often theoretically impressive but unrealizable in practice, as storage subsystems cannot always sustain the throughput. The published info from three leading hyperscalers, shown in Tables 10 should make this aspect more clear.
Table 7. Google Cloud Platform (GCP) virtual instance parameters.
Service Maximum Bandwidth Key Features & Optimizations
Compute Engine VMs (e.g., A3, G4, C4, H4D) Up to 1,000 Gbps (A3-highgpu-8g) and 400 Gbps (G4-standard-384) Tier1 networking is required on selected machine types to reach 100/200 Gbps per-VM egress. Google Virtual NIC (gVNIC) and Fast Socket improve performance for distributed workloads.
Dedicated Interconnect 10 Gbpsand 100 Gbps port speeds Direct, private physical connection from on-premises network to Google for high-throughput hybrid connectivity.
Table 8. AWS virtual instance parameters.
Service Maximum Bandwidth Key Features & Optimizations
EC2 Instances (e.g., HPC, GPU, Network-optimized) Up to 400 Gbps Requires specific high-end instances. Enhanced Networking (ENA) for high packet-per-second, low-latency performance.
AWS DirectConnect 1–100 Gbps (Dedicated) Dedicated network from premises to AWS. Supports Link Aggregation Groups (LAG) and single 100 Gbps connections.
Table 9. Azure virtual instance parameters.
Service Maximum Bandwidth Key Features & Optimizations
Azure VMs(HPC/GPU series) Varies by instance, up to 100 Gbps+ Accelerated Networking reduces latency, maximizes throughput; HPC often uses InfiniBand.
Azure ExpressRoute Up to 100 Gbps(Dedicated) Private connection from infrastructure to Azure; ExpressRoute Premium supports higher route/virtual network limits.
Azure Firewall Premium Up to 100 Gbpsthroughput Handles high-volume network traffic with deep packet inspection enabled.
Focusing only on computing, ignoring data availability, and I/O
In the context of using the cloud for HPC, the role of data I/O is defined by the following premises:
  • Without data, compute is useless.
  • Data is produced predominantly at the edge (e.g., LCLS-II (Thayer and Fang, 2019)).
  • Efficiently moving this data to the cloud remains unsolved.
  • Regardless of computational speed, data transport imposes a fundamental delay.
  • Treating this latency (WSJ Staff, 2025) and (Morales, 2025) deserves scrutiny.
Table 10. Typical performance information published by hyperscalers.
Cloud Provider Workload Metric Throughput Implication
AWS Western Digital HPC Simulation 2.5M tasks on 1M vCPU cluster in 8 hours Massive-scale, low-latency, high-bandwidth communication; reduces job time from 20 days on-prem to 8 hours in cloud.
[…]
Azure Univ. of Bath HPC 250 on-prem nodes $\rightarrow$ thousands in Azure Provides on-demand, high-speed burst capacity; ultra-high data access elasticity.
Lack of reproducibility
Results are presented primarily in marketing form and omit key engineering details required for reproducibility, including storage, network, system configuration.
3.6.3. The Antidote: A Co-Designed Data Path for the Cloud
The solution is not to work within the existing cloud models’ constraints, but to architect a parallel data path coexisting with it. By embedding zx in DPUs (NVIDIA, 2025b; Zettar Inc., 2024), we create a high-performance path that bypasses the cloud software stack. This provides:
  • File Transfer Semantics: Direct file-to-file transfer, avoiding complex HTTP chunking.
  • Line-Rate Performance: Latency-insensitive, encrypted movement at full line rate.
  • Co-existence: Zero changes to REST APIs, improving performance without disrupting services.
This architecture mitigates cloud data movement penalties and equips cloud providers with high-performance data mobility for efficient, dynamic data circulation.
3.6.4. Empirical Evidence of the Cloud Data Movement Limitations
The performance penalties imposed by standard cloud data paths are measurable. Using the GoToCloud platform on AWS (Moriya et al., 2024), a comparative analysis conducted by High Energy Accelerator Research Organization (KEK) (High Energy Accelerator Research Organization (2025), KEK) vividly illustrates this point. As summarized in Fig. 11, the transfer of a 1.2 TiB Cryo-electron microscopy (Cryo-EM) dataset was measured with zx vs. aws-cli
Figure 11. Transfer comparison of a 1.2 TiB Cryo-EM dataset from KEK to AWS Regions using aws-cli and zx.
A bar chart comparing transfer performance of zx vs. aws-cli for Cryo-EM dataset transfers from KEK to AWS Tokyo and Virginia.
Results:
  • Over 63 km (KEK → AWS Tokyo), zx completed in 22.66 minutes.
  • Over 10,851 km trans-Pacific (KEK → AWS N. Virginia), zx completed in 39.97 minutes, a 1.76x increase in transfer time.
This performance differential is notable given the 10 Gbps network at KEK. While partially constrained by a firewall, zx utilized the bandwidth effectively. In comparison, aws-cli transfer to N. Virginia required 235.18 minutes.
Two technical observations:
  1. (1)
    Cloud’s Self-Imposed Bottleneck: The underperformance of aws-cli indicates that the bottleneck is not the physical network, but the inefficient software stack and limited end-to-end control.
  2. (2)
    Architectural Solution: zx demonstrates that long-distance transfers can be latency-insensitive. Embedding this in DPUs would eliminate this performance delta, bringing cloud data movement close to optimized bare-metal transfers.
This case study moves beyond benchmarking to demonstrate a fundamental performance trade-off in the standard cloud model for fast data movement.
4. Methodology and Reproducibility
To enable independent validation of the performance claims presented in this paper, we have provided all essential configuration details required to replicate our testing environment. This includes the exact kernel parameters, latency simulation values, and visualization commands, all of which are publicly available in a GitHub repository (Fang, 2025). The iperf3 benchmark tool used in comparative analysis is also publicly available from its official repositories (man7.org, 2025).
The zx data mover and associated deployment tooling represent proprietary commercial technology reserved for customers and partners. The provided materials allow independent verification of the benchmarking methodology and performance claims without requiring access to the commercial zx software.
5. Deployments
5.1. Validation Platforms
This testing is part of a decade-long, sustained effort to validate the co-design principle across diverse production environments, as documented in prior studies (Thayer and Fang, 2019; Energy Sciences Network (2017), ESnet; Energy Sciences Network (2018), ESnet; ICM, 2019; NSCC, 2019; Fang et al., 2016; Fang, 2019).
Testing for this study was performed on multiple platforms, demonstrating the adaptability of the co-design principle. The validation platforms included:
  • A testbed built using two HPE DL380 Gen 11 servers and two Intel M50CYP2UR208 servers at the former Intel Swindon Lab (U.K.).
  • Two HPE DL380 Gen 11 server-based appliances located in Switzerland and California, U.S. data centers, connected via a shared production 100 Gbps link carrying other network traffic.
  • Two mini appliances deployed at two distinct sites in Taiwan, connected via the Taiwan Advanced Research and Education Network (TWAREN) (TaiWan Advanced Research and Education Network (2025), TWAREN).
  • Independent validation conducted at KEK in Japan using the GoToCloud platform on AWS, as reported in Section 3.6.4.
5.2. Data Availability and Licensing
In the spirit of reproducible research, all elements required to replicate the testing and benchmarking methodology presented in this paper are provided in a public GitHub repository (Fang, 2025). The zx software data mover is a commercially licensed product and is not available for public distribution.
6. Conclusions
The arguments presented in this paper against over-reliance on isolated, formulaic optimizations are rooted in a system perspective. As in the design of composite laminates (Fang and Springer, 1993), where macroscopic properties of a composite plate result from the arrangement of discrete piles, the performance of a data movement system emerges from the interplay of discrete, non-linear components: CPU cores, memory hierarchies, storage devices, network interfaces, among others.
The holistic co-design principle central to this work represents the evolution of a decade-plus sustained and continuing engineering effort. This lineage traces back to foundational architectural blueprints developed at SLAC National Accelerator Laboratory (Fang et al., 2016) and subsequently validated through formal DOE/ESnet evaluation (Kissel and Fang, 2020c). The architecture’s efficacy has been demonstrated through high-stakes production trials at the Linac Coherent Light Source (LCLS) (Thayer and Fang, 2019) and through extensive public discourse (Fang, 2019), (Fang and Kissel, 2021), (Fang, 2021). This maturing perspective has recently been extended to modern hardware platforms, demonstrating that consistent, high-performance data movement can be seamlessly carried out on Data Processing Units (DPUs) (Fang, 2022), (Zettar Inc., 2024). This maturing perspective underscores that consistent, high-performance data movement has always been fundamentally a systems engineering challenge rather than an isolated network optimization problem.
This work also introduces a validated path toward democratizing high-performance data movement. The principle of co-design integrates the data mover software, host OS, and hardware stack, thereby eliminating the reliance on complex, manual, and often proprietary system tuning, resulting in a straightforward and cost-effective deployment model that provides predictable outcomes from the resource-constrained edge sites to the high-throughput core. This consistency is evidenced by the use of commonly available hardware valued at approximately \$2,000 per device for the 1–10 Gbps mini-appliances. This capital efficiency proves that rich data transfer capabilities are no longer confined to environments with multi-million dollar testbeds or dedicated network engineering staff. The resulting architecture is not merely faster, but fundamentally more accessible to the Research and Education communities at the perimeter.
The enduring lesson is that seeking a simple, elegant formula to govern such systems is a mirage. The true path to performance lies not in mathematical purity applied to a simplified model, but in embracing the inherent complexity through empirical, holistic co-design. This co-design is required on two fundamental levels: optimizing the internal host architecture and ensuring efficient, scale-out coordination among peer appliances. This paper demonstrates that the same systems-thinking principles that govern the design of physical materials are equally critical for architecting the high-performance data systems of the digital age.
The scalability of this architectural approach was demonstrated as early as 2018 in the petabyte-transfer setup (Energy Sciences Network (2017), ESnet) and (Energy Sciences Network (2018), ESnet) shown in Fig. 12 (reproduced from (Kissel and Fang, 2020c) Appendix 6.1). This scale-out design used multiple older servers with zx-aggregated 10 Gbps interfaces, proving that the invariant software architecture could distribute workload across nodes without cluster management overhead. This consistency ensures that the logical appliance design—whether implemented as a 10 Gbps mini-appliance or a 400 Gbps core node—remains uniform across scales. For example, modern 2U2N servers, e.g., the SYS- [^2] BT-DNR (Supermicro, 2025) (Fig. 13), enable the same scale-out efficiency within a single chassis. It extends our testbed-based methodology to 400 Gbps and beyond, using the same architectural principles that successfully moved petabytes in 2018.
Figure 12. Zettar’s testbed at SLAC from 2015 to 2019. See page 19 of (NSCC, 2019) for more details of the configuration. (Copyright Zettar Inc. Reproduced with permission for this publication.)
Photograph of Zettar’s testbed at SLAC National Accelerator Laboratory from 2015-2019 showing server rack setup.
Figure 13. The Supermicro 2U 2-Node BigTwin with 12 hot-swap 2.5" NVMe drives per node (Diagram by Chin Fang).
Photograph of Supermicro 2U 2-Node BigTwin server with 12 hot-swap 2.5" NVMe drives per node.
The practical impact of this architectural approach extends beyond technical metrics: predictable, high-throughput data movement enables research, education, and industrial workflows at scales previously reserved for multi-million-dollar infrastructures, translating high-performance capability into tangible strategic and economic value.
7. Acknowledgment
The authors extend their sincere gratitude to the following colleagues for their invaluable insights and for graciously reviewing early drafts of this manuscript:
  • Chin Guok (CTO, ESnet) for reviewing the manuscript and providing network expertise.
  • Dr. Ezra Kissel (Network Research Engineer, ESnet) for his thoughtful review of our testbed methodology and for his foundational 2012 work on Linux network emulation at the University of Delaware.
  • Dr. Amedeo Perazzo (formerly LCLS-II Controls and Data Systems Director, SLAC) for offering insights from large-scale scientific facilities.
  • Dr. Jana Thayer (LCLS Experimental Data Systems Division Director, SLAC) for her review and for insights into the LCLS-II data system architecture and real-time processing requirements that motivated our focus on streaming data movement and holistic system design.
  • Dr. Wilko Kroeger (LCLS-II Information Specialist, SLAC) for his insights and review of the data management challenges associated with the LCLS-II, which informed the architectural perspectives in this work.
  • Mark Gray (Head of Strategic Partnership, and formerly Head of Scientific Platforms, Pawsey Supercomputing Centre) for contributing HPC and storage perspectives, sharpened by our long-standing discussions on SKA data movement challenges since 2018.
  • Sven Breuner (creator of BeeGFS and the elbencho tool) for his review and for insights from a storage architecture and performance perspective.
  • Chih Chuan Shih (Platform Lead, Genome Institute of Singapore) for providing genomics and data lifecycle insights.
  • Dr. Tsukasa Nakamura (KEK, Institute of Materials Structure Science) for his assistance in establishing the KEK IT and internal network environment for the zx implementation, leveraging his expertise in large-scale computational research and bioinformatics.
Their feedback, drawn from deep expertise across networking, large-scale scientific facilities, high-performance computing, high-performance storage, and genomics, significantly strengthened this work. The views and conclusions presented herein are, of course, solely those of the authors.
The authors also gratefully acknowledge the generous in-kind support that enabled critical empirical validation of the architecture: SLAC National Accelerator Laboratory for providing testbed space and facilities (2015–2019); ESnet for provisioning a dedicated 5,000-mile 100 Gbps On-demand Secure Circuits and Advance Reservation (OSCAR) loop (2015–2019); Mellanox Technologies (now part of NVIDIA) and Intel Corporation for providing essential hardware components; and AIC for providing the storage servers.
Chin Fang wishes to thank his colleague Riccardo Veraldi for valuable assistance with TCP congestion control and kernel optimization in 2024. While Chin was building the data movement appliances (Section s33), Riccardo’s guidance contributed to the success of this work.
Chin Fang offers special thanks to Dr. Roger Leslie (Les) Anderton Cottrell, whose mentorship prior to his retirement from SLAC in 2017 played a pivotal role in enabling further collaboration with the DOE national laboratory community.
Chin Fang also wishes to acknowledge two Zettar colleagues, Igor Solovyov and Alexander Nazarenko, for their critical engineering contributions to the zx software, which proved instrumental in the work presented herein. Mr. Nazarenko, paired with Chin Fang, earned the Overall Winner title at the SCA19 Data Mover Challenge (NSCC, 2019).
Finally, Chin Fang wishes to express gratitude to his wife for her unwavering personal support throughout the duration of this sustained effort.